An ideal vapor-compression refrigeration cycle operates at steady state with Refrigerant 134a as the working fluid. Saturated vapor enters the compressor at 1.25 bar, and saturated liquid exits the condenser at 5 bar. The mass flow rate of refrigerant is 8.5 kg/min.

A. Determine the magnitude of the compressor power input required, in kW (report as a positive number).

B. Determine the refrigerating capacity, in tons.

C. Determine the coefficient of performance.

Please answer all parts of the question. Thanks!
Answer #1

Draw the T-s diagram of the refrigeration cych Obtwining data from probenty tabled : T 5 bar 1.25 bay - State 0 :: P2 = 1.25(b)! Refrigerating capacity? - Qin = m (the-ha) = (s. ksls) x (234-45 - 71-33) hehe in = 23.108 kW ton of refregiration = 3-5